OpenStudy (anonymous):

For 2x - 3y = -10, when x = 4, What is the value of y

OpenStudy (whpalmer4):

wherever you see "x" in the equation, write "(4)". then solve for y...

OpenStudy (anonymous):

Substitute x for 4 in your equation, so you have 2(4) - 3y = -10 8 - 3y = -10 8 - 8 - 3y = -10 - 8 -3y = -18 -3y/-3 = -18/-3 y = 6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np